OpenStudy (anonymous):

Someone help me pleaseee Solve the equation (X-3)(x+5)=-7

OpenStudy (anonymous):

x^2+2x-15=-7 x^2+2x-15+(7)=-7(+7)

OpenStudy (anonymous):

x^2+2x-8=0

OpenStudy (kendricklamar2014):

Step 3: Factor left side of equation. (x−2)(x+4)=0 Step 4: Set factors equal to 0. x−2=0 or x+4=0 x=2 or x=−4

OpenStudy (kendricklamar2014):

Answer: x = 2 or x= -4

OpenStudy (kendricklamar2014):

That is the answer @natth88